OpenStudy (anonymous):

Fred is fencing his yard, which is shaped as a rhombus. He used 4x meters of fencing material for one side and (x + 9) meters for the adjacent side. ----------------------------- What is the perimeter of the yard?

OpenStudy (anonymous):

Do I just do 4x = x + 9?

OpenStudy (anonymous):

^^ Nevermind that didn't fit one of the answer choices. lol.

OpenStudy (diyadiya):

is 48 one of your answers ?

OpenStudy (anonymous):

Yep! How'd ya do it?

OpenStudy (diyadiya):

Ok right i'll tell you xD

OpenStudy (diyadiya):

you were right All sides of a rhombus is congruent so 4x=x+9 3x=9 x=3 got it so far ?

OpenStudy (anonymous):

Yeah I do

OpenStudy (diyadiya):

Ok so in the question its given he used 4x meters for fencing one side that means 4x = 4*3 =12m ( since x=3) and x+9 for other side (3)+9 =12m So you have one side of Rhombus as 12m Perimeter = Sum of all sides = 12+12+12+12 (or 4 * 12) =48 m :D

OpenStudy (diyadiya):

Do tell me if you don't understand any part of the answer :)

OpenStudy (anonymous):

oh.. All I had to do was multiply it x_x Thank you haha

OpenStudy (anonymous):

and plug it in*

OpenStudy (diyadiya):

Yess LoL :) You're Welcome !!